1. If you get 11 hours of sleep at night how many seconds of sleep is that? if you get an average of 5 hours of sleep in a year how many days you consumed sleeping Throughout the year?
2. Lance run in an uphill With the distance of 12 km in four hours and walk 80 m in 60 seconds. what is the total distance he covered in meters? how much time he consumed in seconds?
3. Samuel is planning to join a basketball team And one of the requirements is that height must be 5 feet and 9 inches tall, If his heart is 192 cm, Does he qualify in height requirement?
1
Expert's answer
2021-08-27T14:40:03-0400

1)


"t=11(3600)=39600\ s\\"

"N=\frac{5}{24}365=76\ days"


2)


"s=12000+80=12080\ m\\
t=4(3600)+60=14460\ s"


3)


"5(12)+9=69\ in=175\ cm"

Answer: No